# Cyclic Vision-Language Manipulator: Towards Reliable and Fine-Grained Image Interpretation for Automated Report Generation

### v2

@@ 5 Discussion
While there is an abundance of X-ray report generators, existing evaluation metrics primarily focus on assessing the accuracy of the generated language and the detected abnormalities. Our proposed method introduces a novel metric by enabling the verification of identified features, facilitating the comparison of image features used for the generated findings in the report. This provides a unique approach to evaluating the trustworthiness of the generated content from different report generators by examining whether they rely on clinically consistent image features.
5.3 Limitation and Future works
@@
Currently, our method relies on CheXbert, which classifies only 13 diseases for reorganizing the report. This limitation restricts CVLM to explaining these 13 specific keywords. However, given the success of large language models (LLMs), such as GPT-4, in accurately organizing reports, as demonstrated in a recent work Zhang et al. (2024), we can easily extend the proposed method to handle open vocabulary explanations by leveraging LLMs to extract a broader list of abnormalities from the generated reports, thereby enhancing the flexibility and applicability of CVLM. Besides, we will include more human experts and return them the findings from different report generators and their corresponding visual contexts, to further benchmark the reliability of the existing report generators.
@@ B.1 Comparison of different report generators using CVLA
In this section, we present supplementary examples using two query X-rays and the explanation results generated by two different report generators, R2GenCMN and R2GEN. The examples are illustrated in Figures 12 and 13. The counterfactuals generated from these images remove the findings by feeding the regenerated reports back into the abnormality classification models.
@@ B.3 Ablation study in terms of inference times
We analyzed the influence of the denoising step $T$ during diffusion model inference on achieving cyclic success. We evaluated $T$ values ranging from 20 to 100 and found that $T>50$ yields the highest success rates (SRs).

### v3

@@ 4.3.1 Broader Applications
While there is an abundance of X-ray report generators, existing evaluation metrics primarily focus on assessing the accuracy of the generated language and the detected abnormalities. Our proposed method introduces a novel metric by enabling the verification of identified features, facilitating the comparison of image features used for the generated findings in the report. This provides a unique approach to evaluating the trustworthiness of the generated content from different report generators by examining whether they rely on clinically consistent image features. The proposed method is modality-agnostic and can be easily extended to report generators for other imaging modalities, such as MRI and CT scans, to improve the explainability of generated reports.
4.3.2 Limitation and Future Work
@@
Currently, our method relies on CheXbert, which classifies only 13 diseases for reorganizing the report. This limitation restricts CVLM to explaining these 13 specific keywords. However, given the success of large language models (LLMs), such as GPT-4 OpenAI (2023), in accurately organizing reports, as demonstrated in a recent work Zhang et al. (2024), we can easily extend the proposed method to handle open vocabulary explanations by leveraging LLMs to extract a broader list of abnormalities from the generated reports, thereby enhancing the flexibility and applicability of CVLM. Besides, we will include more human experts and return them the findings from different report generators and their corresponding visual contexts, to further benchmark the reliability of the existing report generators.
@@ 5 Conclusion
Acknowledgments
[…]
Contribution Statement
Yingying Fang and Zihao Jin (denoted by on the first page) contributed equally to the conceptualization, methodology, investigation, and writing of this work. Shaojie Guo and Jinda Liu supported the benchmarking and visualization. Zhilin Yue contributed to the methodology and visualization. Yijian Gao, Junzhi Ning, and Zhili contributed to the review and editing of the manuscript. Simon Walsh provided clinical evaluation and review. Guang Yang (the corresponding author, denoted by on the first page) contributed to the methodology, as well as the review and editing of the manuscript.
@@ B.1 Comparison of different report generators using CVLA
In this section, we present supplementary examples using two query X-rays and the explanation results generated by two different report generators, R2GenCMN and R2GEN. The examples are illustrated in Figures 10 and 11. The counterfactuals generated from these images remove the findings by feeding the regenerated reports back into the abnormality classification models.
@@ B.2 Comparison of different interpretation methods
Case 5: Analysis of Cases with Generated Milder Findings
We further investigate CVLM‘s ability in explaining the subtle findings in the generated report, such as the findings with adjective ‘mild’ and ‘borderline’ abnormalities. In Fig. 16, we show two cases of detected ‘mild’ and ‘borderline’ cardiomegaly in the report generated by R2GENCMN. The results show that CVLM can localize subtle features associated with generated findings of minor extent.
I M
(a) case of ‘mild’
I M
(b) case of ‘borderline’
Figure 16: Explanations for varying extents of cardiomegaly (initial image (I) and manipulated result (M)))
@@ B.3 Ablation study in terms of inference times
We analyzed the influence of the denoising step $T$ during diffusion model inference on achieving cyclic success. We evaluated $T$ values ranging from 20 to 100 and found that $T>50$ yields the highest success rates (SRs)in Table 2. We also compared the inference time of the proposed method with RGRG and cross attention methods in Table 3.
@@
Table 3: Comparison between different explanation method in terms of inference times
Methods CVLM (Ours) RGRG Cross Attention
Time (s) 5.16 4.50 5.00